# The Unpaid Toll: Quantifying the Public Health Impact of AI

### v1

@@ 2.2.2 Scope 2
The combustion of fossil fuels for electricity production is a major emitter of criteria air pollutants, releasing large amounts of PM2.5, SO2, NOx, VOCs, and others. [^4] Wet cooling towers, including those used by data centers [9, 10] and carbon-free nuclear power plants, rely on water evaporation for heat rejection and produce PM2.5 due to spray drift droplets [62, 63]. Nonetheless, because of limited data available, we exclude the cooling tower PM2.5 emission from our analysis unless other specified. Critically, the growing energy demands of AI are already delaying the decommissioning of coal-fired power plants and increasing fossil-fuel plants in the U.S. and other countries [6, 40]. For example, in addition to keeping 2,099 MW coal generation capacity until 2039 (more than 80% of the 2024 level), Virginia Electric and Power Company plans to install 5,934 MW gas-fired plants to meet the growing energy demand driven by AI data centers [41]. At the national level, per the EIA’s projection, [39], the 2050 natural gas consumption for U.S. electricity generation will be about 80% of the 2024 level in the baseline reference case, and even exceed the 2024 level by 20% if the zero-carbon technology cost is high; for coal consumption by the electricity sector in 2050, the numbers will also be considerably high, about 30% and over 50% of the 2024 level in the baseline reference case and in the high zero-carbon technology cost case, respectively. These projections were published by the EIA at the very beginning of the generative AI boom in early 2023. More recently, it has been reported that AI data centers could even be primarily powered by coal power plants in some countries [40]. As a result, AI’s scope-2 air pollution is expected to remain at a high level for a substantially long time into the future.
@@ 2.2.3 Scope 3
The surging demand for AI necessitates large quantities of computational hardware, including graphics processing units (GPUs), thus intensifying the supply chain requirements [64]. However, semiconductor manufacturing generates various criteria air pollutants, wastewater, toxic materials, and hazardous air emissions [18]. Moreover, the energy-intensive nature of semiconductor production further contributes to pollutants from power plants. Combined with other pollution sources such as transportation and electronic waste recycling [65], the supply chain activities form a large portion of AI’s scope-3 impact on public health.
@@
Although semiconductor manufacturing facilities are subject to air quality regulations [66], they still pose significant risks, affecting populations across large regions. Maricopa County, AZ, has been an EPA-designated non-attainment area for several years due to its failures to meet federal air quality standards [67]. The establishment of multiple semiconductor facilities in such areas could further exacerbate air quality issues. In 2023–2024, the estimated annual public health impact of a single semiconductor facility was \$26-39 million, with \$14-21 million attributed to direct on-site emissions of air pollutants from the facility, based on COBRA estimates [18, 46]. Moreover, relocating the facility to a planned site in Licking County, Ohio, could nearly quadruple public health costs to \$94-156 million, with \$23-36 million resulting from direct on-site emissions. This increase is partly due to Ohio’s weather conditions and higher reliance on coal-based power [68]. The details of calculations are available in Appendix A.4. Importantly, the global demand for AI chips in 2030 is projected to be tens of times of the overall production capacity of this single facility [69], further magnifying the overall scope-3 public health impact of AI. It is also worth noting that additional pollutants, including hazardous air pollutants like hydrogen fluoride, may further elevate public health costs but are not included in this analysis.
3 Quantifying the Public Health Impact of AI
@@
To quantify the public health impact of AI, we present a general methodology that quantifies AI’s criteria air pollutants at the emission source, models its dispersed air pollutants at different receptors (i.e., destination regions), and finally obtains the public health impact and cost at each receptor.
@@
For an AI task (e.g., AI model training), we consider $M$ types of criteria air pollutants, $N$ receptor regions of interest (e.g., all the U.S. counties), $H$ types of public health impacts (e.g., mortality, asthma symptoms, school loss days, etc.). We use $p^{s}=(p_{1}^{s},\cdots,p_{M}^{s})$ and $p_{i}^{r}=(p_{i,1}^{r},\cdots,p_{i,M}^{r})$ denote the quantities for $M$ types of air pollutants attributed to the task at the emission source and at the receptor $i$, respectively, for $i=1,\cdots,N$. Additionally, we use $h_{i}=(h_{i,1},\cdots,h_{i,H})$ and $c_{i}=(c_{i,1},\cdots,c_{i,H})$ to denote the incidences and economic costs associated with $H$ types of health impacts at receptor $i$, respectively, for $i=1,\cdots,N$. With a slight abuse of notations, we reuse these symbols when modeling AI’s public health impacts across the three different scopes.
@@ 3.1 Criteria Air Pollutants at the Source
We first model AI’s criteria air pollutants at the source across the three different scopes in Section 2.2.
3.1.1 Scope 1
On-site backup diesel generators are sized based on the data center power capacity and routinely tested to ensure a high availability of the entire data center. Thus, the overall scope-1 air pollutants should be attributed to each computing task based on its power allocation and duration. Suppose that the overall scope-1 emission by an AI data center under consideration is $\bar{p}^{s}=(\bar{p}_{1}^{s},\cdots,\bar{p}_{M}^{s})$, for $M$ types of air pollutants, over a timespan of $\overline{T}$ (e.g., one year). Considering an AI task that is allocated a fraction of $x\in(0,1]$ of the overall data center power capacity and lasts for a duration of $T$, we express the scope-1 air pollutants attributed to the AI task as
@@ 3.1.1 Scope 1
3.1.2 Scope 2
AI’s scope-2 air pollutants come from its usage of electricity generated from fossil fuels. Suppose that the power grid serving the AI data center has an emission rate of $\gamma=(\gamma_{1},\cdots,\gamma_{M})$ for $M$ types of air pollutants to produce each unit of electricity. In practice, the power grid consists of multiple interconnected power plants to supply electricity to many customers over a wide area (e.g., a balancing area [70]). Thus, similar to carbon footprint accounting [71], the air pollutant emission rate $\gamma$ can be calculated based on either the weighted average emission rate of all the power plants (i.e., $\gamma=\frac{\sum_{k}\gamma_{k}\cdot b_{k}}{\sum_{k}\cdot b_{k}}$ where $\gamma_{k}$ and $b_{k}$ are the emission rate and electricity generation of the power plant $k$) or the emission rate of the marginal power plant (i.e., the power plant dispatched in response to the next electricity demand increment), which are referred to as average emission rate or marginal emission rate, respectively. The average emission represents a proportional share of the overall air pollutant emission by an electricity consumer, while the marginal emission is useful for quantifying the additionality of air pollutants due to a consumer’s electricity usage.
Suppose that the electricity consumption by the AI task is $e$, including the data center overhead captured by the power usage effectiveness. Then, we can write the scope-2 air pollutants as
@@ 3.1.2 Scope 2
which is either based on either average attribution or marginal attribution. While the marginal emission is typically associated with a single marginal power plant, the average emission is spread across all the interconnected power plants within a wide area such as a power balancing area [71, 70]. Thus, when considering the average attribution method, we split the energy consumption $e$ over all the power plants in proportion to their contributions to the grid’s supply and calculate the corresponding per-plant emission. In other words, each involved power plant is an individual pollution source, and the air pollutant emission at the $k$-th power plant is $p_{k}^{s}=e\cdot\frac{b_{k}}{\sum_{k}b_{k}}\cdot\gamma_{k}$, where $b_{k}$ is the electricity generation of the $k$-th power plant.
Since both the average and marginal air pollutant emission rates vary over time and locations to meet the supply-demand balance, we can also refine the calculation of scope-2 air pollutants in (2) by considering the summation of air pollutants over multiple time slots over the AI task’s duration.
3.1.3 Scope 3
Following the attribution method for scope-3 carbon emission and water consumption [13, 9], we attribute the AI hardware’s air pollutants during the manufacturing process to a specific task based on the task duration. Specifically, let the AI hardware’s expected lifespan be $\overline{T}_{0}$ and the AI task lasts a duration of $T$. Considering that the $M$ types of air pollutants for manufacturing the AI hardware are $\bar{p}_{0}^{s}=(\bar{p}_{0,1}^{s},\cdots,\bar{p}_{0,M}^{s})$ and excluding other miscellaneous pollutants (e.g., transportation), we obtain AI’s scope-3 air pollutants as
@@ 3.1.3 Scope 3
As an AI server cluster includes multiple hardware components (e.g., GPU and CPU) manufactured in different locations, we apply (3) to estimate the scope-3 air pollutants for each component manufactured in a different location.
@@ 3.2 Air Quality Dispersion Modeling
The movement of air pollutants can be modeled using mathematical equations to simulate the atmospheric processes governing the dispersion, known as dispersion modeling. By incorporating emission data and meteorological inputs, dispersion modeling can predict pollutant concentrations at selected receptor locations [72]. We consider a general dispersion model $(p^{r}_{1},\cdots,p^{r}_{N})=D_{\theta}(p^{s})$, which yields the amount of $M$ types of air pollutants $p_{i}^{r}=(p_{i,1}^{r},\cdots,p_{i,M}^{r})$ at the receptor region $i=1,\cdots,N$. The parameter $\theta$ captures the geographical conditions, emission source characteristics (e.g., height), and meteorological data if applicable [73]. We apply the dispersion model to each scope of air pollutants (Section 3.1) to estimate the corresponding pollutant concentrations at receptor regions.
In practice, many dispersion modeling tools are available, including AERMOD, CTDMPLUS, PCAPS and InMAP with a reduced complexity [72, 22, 74, 75]. For example, PCAPS (Pattern Constructed Air Pollution Surfaces), an advanced reduced-complexity model that provides representations of both primarily emitted PM2.5 and secondarily formed PM2.5 and ozone, is used in COBRA as a quick assessment of otherwise lengthy iterations and simulations of various pollution scenarios in terms of the annual average PM2.5 and seasonal average maximum daily average 8-hour ozone [75, 22]. Even compared with state-of-the-science photochemical grid models, PCAPS provides similar prediction accuracies and can realistically capture the change in air pollution due to changing emissions [75]. More specifically, for electric power sectors and on-road/highway vehicle sectors (the two sectors we consider in Section 4), the prediction results of PCAPS compare very well with photochemical model predictions, with Pearson correlation coefficients of 0.92 and 0.94, respectively [22, 75].
@@ 3.3 Converting Health Outcomes to Economic Costs
By assessing pollutant levels $p_{i}^{r}=(p_{i,1}^{r},\cdots,p_{i,M}^{r})$ and population size at each receptor region $i$, we can estimate the incidences of health outcomes $h_{i}=(h_{i,1},\cdots,h_{i,H})$ and the corresponding public health cost $c_{i}=(c_{i,1},\cdots,c_{i,H})$. The relations between $p_{i}^{r}$ and $h_{i}$ and between $h_{i}$ and $c_{i}$ can be established based on epidemiology research [22]. For example, the premature mortality rate can be modeled as a log-linear function in terms of the PM2.5 level [23].
@@
3.4 Implementation
We now briefly describe the specific implementation we use to study the public health impact of U.S. data centers and AI training. The details are available in Appendix A.
Due to the limited data available for scope-1 and scope-3 impacts, we mainly focus on the scope-2 health impacts from electricity consumption. To account for future uncertainties, we use the U.S. data center electricity consumption data provided by EPRI [5] and McKinsey [4] under various growth-rate scenarios, excluding cryptocurrency servers. Unless otherwise specified, we consider the average attribution method by default, i.e., attribute the overall health impact within an electricity region to data centers in proportion to their electricity consumption.
To model the air pollutant dispersion and quantify health impacts, we use the latest COBRA (Desktop v5.1, as of October 2024) provided by the U.S. EPA [46]. COBRA integrates reduced-complexity air dispersion modeling (including both primarily emitted PM2.5 and secondly formed PM2.5 and ozone [75]) with various concentration-response functions [22], offering a quantitative screening analysis particularly suitable for large-scale health impacts. The same or similar reduced-complexity modeling tools have been commonly used in the literature to examine the health impacts of various industries over a large area [76, 74], including electric vehicles [77], bitcoin mining [78], and inter-region electricity imports [79], among others. While each health impact model used by COBRA considers 95% confidence intervals, the high-end and low-end estimates provided by COBRA are based on different models instead of the 95% confidence interval of a single model [22]. COBRA provides data for county-level population, health incidence, and valuation projections in 2030, but the baseline emissions are missing [46]. Thus, to account for model uncertainties, we estimate the 2030 baseline emission by extrapolating the COBRA data for 2016, 2023, and 2028 using three extrapolation methods (Linear, Exponential, and Unchanged) as detailed in Appendix A.1.
We only consider the contiguous U.S. and simply refer to it as the U.S. For consistency with COBRA, cities considered county-equivalents for census purposes are also referred to as “counties” in our paper. All our monetary values are for one year (or one AI task if applicable) and in 2023 U.S. dollars.

### v2

@@ 3.2 Scope 2: Power Plants
Along with transportation and manufacturing, the combustion of fossil fuels for electricity production is a leading anthropogenic source of criteria air pollutants, releasing large amounts of PM2.5, SO2, NOx, VOCs, and others [30]. [^5] Wet cooling towers, including those used by data centers [9, 10] and carbon-free nuclear power plants, rely on water evaporation for heat rejection and produce PM2.5 due to spray drift droplets [63, 64]. Nonetheless, because of limited data available, we exclude the cooling tower PM2.5 emission from our analysis unless other specified. More alarmingly, the growing energy demands of AI data centers are already delaying the decommissioning of coal-fired power plants and driving the expansion of fossil-fuel-based plants in the U.S. and globally [6, 65, 66]. For example, in addition to keeping 2,099 MW coal generation capacity until 2039 (more than 80% of the 2024 level), Virginia Electric and Power Company plans to install 5,934 MW gas-fired plants to meet the growing energy demand driven by AI data centers [66].
Figure 3: Public health costs of electricity generation and on-road emissions in the contiguous U.S. in 2023 and 2028 [39]. The error bars represent high and low estimates returned by COBRA using two different exposure-response functions.
Based on the emission data projected by the U.S. EPA’s COBRA modeling tool [39], we show in Fig. 3 that the electric power sector’s total public health cost in the contiguous U.S. is on track to rise, rivaling that of on-road vehicle emissions by all the registered vehicles (including tailpipe exhausts and brakes) in 2028.
Looking forward, the U.S. Energy Information Administration (EIA) projects that coal consumption by the electricity sector in 2050 will still be approximately 30% of the 2024 level if power plants continue operating under rules existing prior to April 2024 [67]. At the global scale, the reliance on coal and other fossil fuels for electricity production has seen little change over the past four decades, highlighting the enduring challenges to fully adopt clean energy for powering data centers [68]. As a result, data centers’ scope-2 air pollution is expected to remain at a high level for a substantially long time into the future.
Although technology companies have started implementing various initiatives—such as purchasing renewable energy credits and nuclear power from small modular reactors [5, 10, 69]—to lower their (market-based) carbon emissions, the vast majority of U.S. data centers remain physically and directly powered by local power grids with a substantial portion of fossil fuel-based energy sources [10]. While the increasingly stringent transmission line constraint is also driving the co-location of data centers with power plants to speed up the construction process, the onsite plants are often gas-fired [70, 71], raising health concerns [31].
@@ 3.3 Scope 3: Supply Chain
The surging demand for AI data centers necessitates large quantities of computational hardware, including graphics processing units (GPUs), thus intensifying the supply chain requirements [72]. However, semiconductor manufacturing generates various criteria air pollutants, wastewater, toxic materials, and hazardous air emissions [18]. Moreover, the energy-intensive nature of semiconductor production further contributes to pollutants from power plants. Combined with other pollution sources such as transportation and electronic waste recycling [73], the supply chain activities form a large portion of data centers’ scope-3 impact on public health.
@@
Although semiconductor manufacturing facilities are subject to air quality regulations [74], they still pose significant risks, affecting populations across large regions. Importantly, the global demand for AI chips in 2030 is projected to be tens of times of the overall production capacity of this single facility [75], further magnifying the overall scope-3 public health impact of data centers. It is also worth noting that additional pollutants, including hazardous air pollutants like hydrogen fluoride, may further elevate public health costs but are not included in this paper.
However, public data on scope-3 criteria air pollutant emissions from semiconductor manufacturers is limited. Thus, we focus on scope-1 and scope-2 emissions, excluding scope-3 impacts from the main analysis. A detailed assessment of the health impacts associated with a specific U.S. semiconductor manufacturing facility is provided in Appendix A.3.
4 Quantifying Task-Specific Public Health Impact
@@
To quantify the public health impact of a specific computing task, we present a principled end-to-end methodology illustrated in Fig. 1. Specifically, the process includes: (1) Quantifying the task’s criteria air pollutants at the emission source; (2) Modeling the dispersed air pollutants at different receptors (i.e., destination regions); (3) Calculating the public health impact and assigning a corresponding monetary value at each receptor.
@@
For a computing task under consideration (e.g., AI model training), we consider $M$ types of criteria air pollutants, $N$ receptor regions of interest (e.g., all the U.S. counties), $H$ types of public health impacts (e.g., mortality, asthma symptoms, school loss days, etc.). We use $p^{s}=(p_{1}^{s},\cdots,p_{M}^{s})$ and $p_{i}^{r}=(p_{i,1}^{r},\cdots,p_{i,M}^{r})$ denote the quantities for $M$ types of air pollutants attributed to the task at the emission source and at the receptor $i$, respectively, for $i=1,\cdots,N$. Additionally, we use $h_{i}=(h_{i,1},\cdots,h_{i,H})$ and $c_{i}=(c_{i,1},\cdots,c_{i,H})$ to denote the incidences and economic costs associated with $H$ types of health impacts at receptor $i$, respectively, for $i=1,\cdots,N$. With a slight abuse of notations, we reuse these symbols when modeling AI’s public health impacts across the three different scopes.
@@ 4.1 Criteria Air Pollutants at the Source
We first model a computing task’s criteria air pollutants at the source across the three different scopes in Section 3.
4.1.1 Scope 1.
Onsite diesel generators are sized based on the data center power capacity, routinely tested to ensure a high availability of the entire data center, and used for demand response. Thus, the overall scope-1 air pollutants should be attributed to each computing task based on its power allocation and duration. Suppose that the overall scope-1 emission by an AI data center under consideration is $\bar{p}^{s}=(\bar{p}_{1}^{s},\cdots,\bar{p}_{M}^{s})$, for $M$ types of air pollutants, over a timespan of $\overline{T}$ (e.g., one year). Considering a task that is allocated a fraction of $x\in(0,1]$ of the overall data center power capacity and lasts for a duration of $T$, we express the scope-1 air pollutants attributed to the task as
@@
4.1.2 Scope 2.
A computing task’s scope-2 air pollutants come from its usage of electricity generated from fossil fuels. Suppose that the power grid serving the data center has an emission rate of $\gamma=(\gamma_{1},\cdots,\gamma_{M})$ for $M$ types of air pollutants to produce each unit of electricity. In practice, the power grid consists of multiple interconnected power plants to supply electricity to many customers over a wide area (e.g., a balancing area [76]). Thus, similar to carbon footprint accounting [77], the air pollutant emission rate $\gamma$ can be calculated based on either the weighted average emission rate of all the power plants (i.e., $\gamma=\frac{\sum_{k}\gamma_{k}\cdot b_{k}}{\sum_{k}\cdot b_{k}}$ where $\gamma_{k}$ and $b_{k}$ are the emission rate and electricity generation of the power plant $k$) or the emission rate of the marginal power plant (i.e., the power plant dispatched in response to the next electricity demand increment), which are referred to as average emission rate or marginal emission rate, respectively. The average emission represents a proportional share of the overall air pollutant emission by an electricity consumer, while the marginal emission is useful for quantifying the additionality of air pollutants due to a consumer’s electricity usage. Suppose that the electricity consumption by the computing task is $e$, including the data center overhead captured by the power usage effectiveness. Then, we write the scope-2 air pollutants as
@@
which is either based on either average or marginal accounting.
To evaluate the public health impacts of U.S. data centers, we consider the average attribution method unless otherwise noted, which is also the standard methodology of carbon emission accounting used by technology companies in their sustainability reports [10, 78, 37].
We can also refine the calculation of scope-2 air pollutants in (2) by considering the summation of air pollutants over multiple time slots over the task’s duration.
Location-based emission. There are two types of scope-2 carbon emission accounting associated with electricity consumption: location-based and market-based [10]. Specifically, location-based carbon emissions refer to the physical carbon emissions attributed to an electricity consumer connected to the power grid, while market-based carbon emissions are net emissions after applying reductions due to contractual arrangements and other credits (e.g., renewable energy credits). As noted by a recent study on carbon accounting [79], location-based accounting is considered essential, whereas market-based accounting is valuable. Moreover, market-based accounting relies on market instruments whose detailed information is often not publicly disclosed. Thus, in this paper, we follow the literature [8] and focus on location-based accounting for scope-2 criteria air pollutants without considering market-based pollution reduction mechanisms.
We also note that market-based emission reduction is likely less effective to mitigate the public health impact. The reason is that, unlike carbon emissions that have a similar effect on climate change regardless of the emission locations, the public health impact of criteria air pollutants heavily depends on the location of the emission source. For example, the public health impact of using pollutant-intensive electricity generated from a populated region may not be effectively mitigated by the clean energy credits generated elsewhere.
4.1.3 Scope 3.
While our empirical analysis focuses on scope-1 and scope-2 health impacts, we present the scope-3 pollutant attribution method to provide a more complete view. Specifically, following the attribution method for scope-3 carbon emission and water consumption [13, 9], we attribute the computing hardware’s air pollutants during the manufacturing process to a specific task based on the task duration. Specifically, let the hardware’s expected lifespan be $\overline{T}_{0}$ and the task lasts a duration of $T$. Considering that the $M$ types of air pollutants for manufacturing the hardware are $\bar{p}_{0}^{s}=(\bar{p}_{0,1}^{s},\cdots,\bar{p}_{0,M}^{s})$ and excluding other miscellaneous pollutants (e.g., transportation), we obtain the task’s scope-3 air pollutants as
@@
As a server cluster includes multiple hardware components (e.g., GPU and CPU) manufactured in different locations, we apply (3) to estimate the scope-3 air pollutants for each component manufactured in a different location.
@@ 4.2 Air Quality Dispersion Modeling
In practice, dispersion modeling tools are used to track the movement of air pollutants. These tools employ complex mathematical equations to simulate the atmospheric processes governing the dispersion. By incorporating emission data and meteorological inputs, dispersion modeling can predict pollutant concentrations at selected receptor locations [80]. We consider a general dispersion modeling tool $(p^{r}_{1},\cdots,p^{r}_{N})=D_{\theta}(p^{s})$, which yields the amount of $M$ types of air pollutants $p_{i}^{r}=(p_{i,1}^{r},\cdots,p_{i,M}^{r})$ at the receptor region $i=1,\cdots,N$. The parameter $\theta$ captures the geographical conditions, emission source characteristics (e.g., height), and meteorological data [81]. We apply the dispersion model to each scope of air pollutants (Section 4.1) to estimate the corresponding pollutant concentrations at receptor regions.
Several dispersion modeling tools are available, including AERMOD, PCAPS and InMAP with a reduced complexity [80, 22, 82, 83]. For example, PCAPS (Pattern Constructed Air Pollution Surfaces), an advanced reduced-complexity model that provides representations of both primarily emitted PM2.5 and secondarily formed PM2.5 and ozone, is used in COBRA as a quick assessment of otherwise lengthy iterations and simulations of various pollution scenarios in terms of the annual average PM2.5 and seasonal average maximum daily average 8-hour ozone [83, 22]. Even compared with state-of-the-science photochemical grid models, PCAPS provides similar prediction accuracies and can realistically capture the change in air pollution due to changing emissions [83]. More specifically, for electric power sectors and on-road/highway vehicle sectors (the two sectors we consider in Section 5), the prediction results of PCAPS compare very well with photochemical model predictions, with Pearson correlation coefficients of 0.92 and 0.94, respectively [22, 83].
@@ 4.3 Converting Health Outcomes to Economic Costs
By assessing pollutant levels $p_{i}^{r}=(p_{i,1}^{r},\cdots,p_{i,M}^{r})$ and population size at each receptor region $i$, we can estimate the incidences of health outcomes $h_{i}=(h_{i,1},\cdots,h_{i,H})$ and the corresponding public health cost $c_{i}=(c_{i,1},\cdots,c_{i,H})$. The relations between $p_{i}^{r}$ and $h_{i}$ and between $h_{i}$ and $c_{i}$ is captured by an exposure-response function and can be established based on epidemiology research [22]. For example, the premature mortality rate can be modeled as a log-linear function in terms of the PM2.5 level [84].
@@
4.4 End-to-End Modeling
Following the end-to-end process shown in Fig. 1, we now briefly describe our modeling methodology to study the public health impact of U.S. data centers and AI training. The details are available in Appendix A.
To quantify data centers’ scope-1 and scope-2 air pollutant emissions, we use air quality permit data for onsite generators [60] and electricity consumption data, including both historical records and 2028 projections, from the Lawrence Berkeley National Laboratory report [4].
To model the air pollutant dispersion and quantify health impacts, we use the latest COBRA (Desktop v5.1, as of October 2024) provided by the U.S. EPA [39]. COBRA integrates reduced-complexity air dispersion modeling (including both primarily emitted PM2.5 and secondly formed PM2.5 and ozone [83]) with various concentration-response functions [22], offering a quantitative screening analysis particularly suitable for large-scale health impacts. The same or similar reduced-complexity modeling tools have been commonly used in the literature to examine the health impacts of various industries over a large area [85, 82], including electric vehicles [86], bitcoin mining [87], and inter-region electricity imports [88], among others. While each health impact model used by COBRA considers 95% confidence intervals, the high-end and low-end estimates provided by COBRA are based on different models instead of the 95% confidence interval of a single model [22].